OpenStudy (anonymous):

Find the derivative of f(x) = -9 / x at x = 6.

OpenStudy (anonymous):

Remember that a variable in the denominator x^n can be rewritten as x^-n and then the power rule can be applied.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (jhannybean):

\[\large f(x) = -\frac9x \]\[\large f(x) = -9x^{-1}\]\[\large f'(x) =-9(-1)x^{-2}\]\[\large f'(x) =9x^{-2}\]rewrite with positive exponents. \[\large f'(x)= \frac{9}{x^2}\]

OpenStudy (jhannybean):

Now evaluate this function when x= 6.\[\large f'(6)= \frac{9}{(6)^2}= \frac{9}{36} = ?\]

OpenStudy (anonymous):

Oh, I understand.

OpenStudy (jhannybean):

Awesome :)